OpenStudy (anonymous):

is the derivative of 1/1+x^2 = 1=x^2-2x/(1+x^2)^2 ??

OpenStudy (anonymous):

= 1+x^2-2x/(1+x^2)^2

OpenStudy (anonymous):

or is the dervivate of 1 = 0?

OpenStudy (phi):

is the problem \[ \frac{1}{(1+x^2)} \]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how do you do the squared like that?

OpenStudy (anonymous):

i think its -2x/(1+x^2)^2

OpenStudy (phi):

if so, you could write it like this \[ (1+x^2)^{-1} \] this is in the form u^n and \[ \frac{d}{dx} u^n = n\ u^{n-1} \frac{du}{dx} \]

OpenStudy (phi):

how do you do the squared like that? what do you mean? how to type it in ? use the equation editor.

OpenStudy (anonymous):

\[x^2\] wow, i i had no idead that was there, lol

OpenStudy (phi):

i think its -2x/(1+x^2)^2 yes, that is correct. with u= 1+x^2 we first do -1* (1+x^2) ^ (-1 -1) then we do du/dx = d/dx (1+x^2) = d/dx 1 + d/dx x^2 = 0+2x = 2x putting it together -2x/(1+x^2)^2

OpenStudy (anonymous):

thank you, thats what i thought it was, just making sure!

OpenStudy (phi):

"\[" x^2 "\]" without the quotes " will also do the job. if you see an expression that you want to know how to type in right click on the equation and you get a menu. Select "show math as", then select Tex commands. you will see the tex expression that creates it. you can cut and paste it into the input window put it between \ [ and \ ] (no spaces between slash and brackets)